OpenStudy (kylav210):

An aquarium sells cubic fish tanks of different sizes. The length of the small size fish tank is 3 feet. The dimensions of the jumbo size fish tank are double the dimensions of the small size fish tank. Which expression can be used to find the ratio of the volume of the small size fish tank to the volume of the jumbo size fish tank? The whole cube of 6 by 3. The whole square of 3 by 6. The whole cube of 3 by 6. The whole square of 6 by 3.

OpenStudy (anonymous):

Volume of Cube is given by : \(\large a^3\) So: Volume of Smaller Cube is : \(3^3\) Now, length of Larger Cube is : Double of Smaller one : \(2 \times 3 = 6\) So: Volume of Large Cube = \(6^3\) So take the ratio now.. \[\large \frac{3^3}{6^3} \implies (\frac{3}{6})^3\]

OpenStudy (kylav210):

Ooh okay! Great explaining! Thanks(: